OpenStudy (anonymous):

simplify, (3c)squared (3c)cubed

Directrix (directrix):

(3c)squared (3c)cubed indicates the product of [(3c)* (3c)] * [(3c)* (3c)*(3c)] = 3^5 times c^5 = (243 c ^ 5).